The sum of two numbers is 4 times the first.Their sum is 50. Find the numbers.
Found 2 solutions by nyc_function, ewatrrr:
Answer by nyc_function(2741)

The sum of two numbers is 4 times the first.Their sum is 50. Find the numbers.

Let x = first number

Let y = second number

x + y = 4x

x + y = 50

Can you finish?

Answer by ewatrrr(24785)

Hi,
Write as You Read: Let x and y be the numbers
x + y = 4*x
y = 3x
.
Sum of the numbers is 50
x + 3x = 50
4x = 50
x = 12 1/2
y = 37 1/2
.
check your answer
12 1/2 + 37 1/2 = 50